# JOLT-SQL: Joint Loss Tuning of Text-to-SQL with Confusion-aware Noisy Schema Sampling

### v1

@@ Abstract
Text-to-SQL, which maps natural language to SQL queries, has benefited greatly from recent advances in Large Language Models (LLMs). While LLMs offer various paradigms for this task, including prompting and supervised fine-tuning (SFT), SFT approaches still face challenges such as complex multi-stage pipelines and poor robustness to noisy schema information. To address these limitations, we present JOLT-SQL, a streamlined single-stage SFT framework that jointly optimizes schema linking and SQL generation via a unified loss. JOLT-SQL employs discriminative schema linking, enhanced by local bidirectional attention, alongside a confusion-aware noisy schema sampling strategy with selective attention to improve robustness under noisy schema conditions. Experiments on the Spider and BIRD benchmarks demonstrate that JOLT-SQL achieves state-of-the-art execution accuracy among comparable-size open-source models, while significantly improving both training and inference efficiency. Our codes are available at this https URL.
@@
Jinwang Song1, Hongying Zan1††thanks: Corresponding author., Kunli Zhang1, Lingling Mu1, Yingjie Han1, Haobo Hua2, Min Peng3, 1Zhengzhou University, 2Zhengzhou University of Aeronautics, 3Wuhan University, jwsong@gs.zzu.edu.cn, {iehyzan, ieklzhang, iellmu, ieyjhan}@zzu.edu.cn huahbo@zua.edu.cn, pengm@whu.edu.cn
@@ 1 Introduction
Text-to-SQL technology, which transforms natural language questions into executable SQL queries (Deng et al., 2022), aims to break down the technical barriers for users interacting with complex databases, enabling non-professionals to conveniently access desired data. With the widespread adoption of data-driven applications across various domains, the research value and application prospects of Text-to-SQL have become increasingly significant. In recent years, the rise of Large Language Models (LLMs) has injected new vitality into this field, and LLM-based Text-to-SQL methods have made remarkable progress (Hong et al., 2024; Shi et al., 2024).
@@
In this work, we propose a novel method: JOint Loss Tuning of Text-to-SQL with Confusion-aware Noisy Schema Sampling (JOLT-SQL). Our objective is to address two primary issues: (1) Conventional pipeline fine-tuning strategies or multi-task learning approaches often require fine-tuning multiple distinct models or incur increased training time due to the expanded multi-task dataset, leading to higher temporal and hardware costs. (2) During actual inference, the schema information provided by the schema linking module is often imperfect, potentially containing irrelevant items or omitting necessary ones. Existing SQL generation models, when fine-tuned solely on ground truth schema or with random sampled noisy schema, struggle to adequately adapt to such imperfect inputs during inference, thereby affecting SQL generation accuracy.
@@
We further introduce Confusion-aware Noisy Schema Sampling, enhancing model robustness against redundant schema linking results. This technique strategically guides the SQL generation task’s attention by incorporating noisy schema items selected based on the model’s points of predictive confusion. Experimental results show that JOLT-SQL, using the Qwen-2.5-Coder-14B model, achieves 88.4%/88.9% execution accuracy on the Spider Dev/Test set and 64.9% on the BIRD Dev set.
@@ 4.1 Setup
To evaluate the JOLT-SQL method, we conducted experiments on two widely used datasets: Spider (Yu et al., 2018) and BIRD (Li et al., 2023a). Spider comprises 200 databases with multiple tables, and its Training/Dev/Test sets contain 7,000/1,034/2,147 question-SQL pairs, respectively. BIRD contains 12,751 unique question-SQL pairs, 95 large databases with a total size of 33.4 GB, and its Training/Dev/Test sets include 9,428/1,534/1,789 pairs, respectively. Given that BIRD’s Test set is not public, we primarily report results on its Dev set.
@@
In our experiments, we compare recent LLM-based Text-to-SQL baselines in two categories. The Prompting methods include DIN-SQL (Pourreza and Rafiei, 2023), DAIL-SQL (Gao et al., 2024a), CHESS (Talaei et al., 2024), E-SQL (Caferoğlu and Ulusoy, 2024), MCS-SQL (Lee et al., 2025), and MAC-SQL (Wang et al., 2025). The SFT-based methods comprise CodeS (Li et al., 2024), SENSE (Yang et al., 2024b), DTS-SQL (Pourreza and Rafiei, 2024), ExSL (Glass et al., 2025), ROUTE (Qin et al., 2024), DB-Explore (Ma et al., 2025), and BASE-SQL (Sheng et al., 2025). These baselines cover a range of prompting and fine-tuning strategies for Text-to-SQL.
@@ 4.4 Ablation Study
The ablation study confirmed that all contribute to JOLT-SQL’s performance. LBA was vital for capturing complex schema relationships, and its absence markedly impacted results. NSS proved key for improving robustness against noisy schema during inference by promoting adaptation in training, while SA helped align training with inference conditions by focusing attention on relevant schema subsets instead of the full schema. The individual or combined removal of these mechanisms consistently reduced performance. Ultimately, disabling all three components substantially degraded overall performance.
@@ A.1 Comparison with Pipeline Fine-tuning Strategies & Efficiency Analysis
In terms of training cost, the results show that JOLT-SQL requires 5 hours and 5 minutes to complete 3 epochs of training. This duration includes an approximate 8.5% overhead from enabling Confusion-aware NSS compared to its own baseline of 4 hours and 41 minutes (JOLT-SQL without Confusion-aware NSS), a time comparable to the standard SFT approach. Even so, JOLT-SQL is significantly more efficient than DTS-SQL. DTS-SQL, requiring independent fine-tuning for two stages, has a total training time of 7 hours and 10 minutes, which is 52.9% longer than the JOLT-SQL baseline (without Confusion-aware NSS), highlighting the efficiency of JOLT-SQL’s single-stage joint optimization.
@@ B.2 Implementation Details
We consistently used the AdamW8bit optimizer (Dettmers et al., 2022), a cosine annealing learning rate scheduler, and enabled gradient checkpointing. We also utilized the Liger-Kernel (Hsu et al., 2024) for further training efficiency optimization and enabled bfloat16 mixed-precision training. Other training hyperparameters are provided in Table 8. For SQL generation during inference, we employed greedy decoding (temperature=0).

### v3

@@ Abstract
Text-to-SQL, which maps natural language to SQL queries, has benefited greatly from recent advances in Large Language Models (LLMs). While LLMs offer various paradigms for this task, including prompting and supervised fine-tuning (SFT), SFT approaches still face challenges such as complex multi-stage pipelines and poor robustness to noisy schema information. To address these limitations, we present JOLT-SQL, a streamlined single-stage SFT framework that jointly optimizes schema linking and SQL generation via a unified loss. JOLT-SQL employs discriminative schema linking, enhanced by local bidirectional attention, alongside a confusion-aware noisy schema sampling strategy with selective attention to improve robustness under noisy schema conditions. Experiments on the Spider and BIRD benchmarks demonstrate that JOLT-SQL achieves state-of-the-art execution accuracy among comparable-size open-source models, while significantly improving both training and inference efficiency. Our code is available at https://github.com/Songjw133/JOLT-SQL.
@@
Jinwang Song1, Hongying Zan1††thanks: Corresponding author., Kunli Zhang1, Lingling Mu1, Yingjie Han1, Haobo Hua2, Min Peng3, 1Zhengzhou University, 2Zhengzhou University of Aeronautics, 3Wuhan University, jwsong@gs.zzu.edu.cn, iehyzan@zzu.edu.cn
@@ 1 Introduction
Text-to-SQL technology, which transforms natural language questions into executable SQL queries (Deng et al., 2022), aims to break down the technical barriers for users interacting with complex databases, enabling non-professionals to conveniently access desired data. With the widespread adoption of data-driven applications across various domains, the research value and application prospects of Text-to-SQL have become increasingly significant. In recent years, the rise of Large Language Models (LLMs) has injected new vitality into this field, and LLM-based Text-to-SQL methods have made remarkable progress (Hong et al., 2024; Shi et al., 2025).
@@
In this work, we propose a novel method: JOint Loss Tuning of Text-to-SQL with Confusion-aware Noisy Schema Sampling (JOLT-SQL). Our objective is to address two primary issues: (1) Conventional pipeline fine-tuning strategies or multi-task learning approaches often require fine-tuning multiple distinct models or incur increased training time due to the expanded multi-task dataset, leading to higher temporal and hardware costs. (2) During actual inference, the schema information provided by the schema linking module is often imperfect, potentially containing irrelevant items or omitting necessary ones. Existing SQL generation models, when fine-tuned solely on ground truth schema or with random sampled noisy schema, struggle to adapt to such imperfect inputs during inference, thereby affecting SQL generation accuracy.
@@
We further introduce Confusion-aware Noisy Schema Sampling, enhancing model robustness against redundant schema linking results. This technique strategically guides the SQL generation task’s attention by incorporating noisy schema items selected based on the model’s points of predictive confusion. Experimental results show that JOLT-SQL, using the Qwen2.5-Coder-14B model, achieves 88.4%/88.9% execution accuracy on the Spider Dev/Test set and 64.9% on the BIRD Dev set.
@@ 4.1 Setup
To evaluate the JOLT-SQL method, we conducted experiments on two widely used datasets: Spider (Yu et al., 2018) and BIRD (Li et al., 2023a). Spider comprises 200 databases with multiple tables, and its Training/Dev/Test sets contain 7,000/1,034/2,147 question-SQL pairs, respectively. BIRD contains 12,751 unique question-SQL pairs, 95 large databases with a total size of 33.4 GB, and its Training/Dev/Test sets include 9,428/1,534/1,789 pairs, respectively. For these two benchmarks, our models are fine-tuned exclusively on their respective Training sets and evaluated on the corresponding Dev and Test sets.
@@
In our experiments, we compare recent LLM-based Text-to-SQL baselines in two categories. The Prompting methods include DIN-SQL (Pourreza and Rafiei, 2023), DAIL-SQL (Gao et al., 2024a), CHESS (Talaei et al., 2024), E-SQL (Caferoğlu and Ulusoy, 2024), MCS-SQL (Lee et al., 2025), RSL-SQL (Cao et al., 2024) and MAC-SQL (Wang et al., 2025). The SFT-based methods comprise CodeS (Li et al., 2024), SENSE (Yang et al., 2024b), DTS-SQL (Pourreza and Rafiei, 2024), ExSL (Glass et al., 2025), ROUTE (Qin et al., 2024), DB-Explore (Ma et al., 2025), and BASE-SQL (Sheng et al., 2025). These baselines cover a range of prompting and fine-tuning strategies for Text-to-SQL.
@@ 4.4 Ablation Study
The results of the ablation study, detailed in Table 3, confirm that all three core components contribute to JOLT-SQL’s performance. LBA was vital for capturing complex schema relationships, and its absence markedly impacted results. NSS proved key for improving robustness against noisy schema during inference by promoting adaptation in training, while SA helped align training with inference conditions by focusing attention on relevant schema subsets instead of the full schema. The individual or combined removal of these mechanisms consistently reduced performance. Ultimately, disabling all three components substantially degraded overall performance.
@@ A.1 Comparison with Pipeline Fine-tuning Strategies & Efficiency Analysis
In terms of training cost, the results show that JOLT-SQL requires 5 hours and 5 minutes to complete 3 epochs of training. This duration includes an approximate 8.5% overhead from enabling Confusion-aware NSS compared to its own baseline of 4 hours and 41 minutes (JOLT-SQL without Confusion-aware NSS), a time comparable to the standard SFT approach. Even so, JOLT-SQL is significantly more efficient than DTS-SQL. DTS-SQL, requiring independent fine-tuning for two stages, has a total training time of 7 hours and 10 minutes, which is 52.9% longer than the JOLT-SQL baseline (without Confusion-aware NSS), highlighting the efficiency of JOLT-SQL’s single-stage joint optimization. This is because JOLT-SQL learns both tasks from a single input sequence, fundamentally halving the total number of forward/backward passes compared to two-stage methods that process separate data for each task.
@@
A.2 Cross-Dataset Generalization
| Methods | Spider-Syn | Spider-Realistic | Spider-DK | KaggleDBQA |
| JOLT-SQL + Qwen2.5-Coder-7B | 77.2 | 83.0 | 75.9 | 47.0 |
| JOLT-SQL + Qwen2.5-Coder-14B | 79.6 | 84.1 | 79.1 | 52.4 |
Table 8 presents JOLT-SQL’s robustness and generalization capabilities across various challenging datasets. We evaluated its performance on Spider-Syn (Gan et al., 2021a), Spider-Realistic (Deng et al., 2021), Spider-DK (Gan et al., 2021b), and the cross-domain KaggleDBQA dataset (Lee et al., 2021). The results indicate that JOLT-SQL maintains strong performance. Furthermore, it demonstrates solid generalization to entirely unseen databases and query types in KaggleDBQA.
A.3 Performance in Low-Resource Settings
| Training data used |
| 5% | 81.9 | 54.0 | ||||
| 10% | 83.8 | 55.3 | ||||
| 25% | 85.4 | 58.5 | ||||
| 50% | 86.4 | 59.7 | ||||
| 100% | 87.0 | 60.4 |
Table 9 evaluates JOLT-SQL’s performance under low-resource training conditions. We fine-tuned the Qwen2.5-Coder-7B model using various proportions of the Spider and BIRD training set. The results indicate that the model’s execution accuracy degrades gracefully as training data decreases, without a sharp drop. Even with only 5% or 10% of the data, JOLT-SQL still achieves acceptable performance, demonstrating its robustness and practical applicability in scenarios with limited annotated data.
@@ B.2 Implementation Details
We consistently used the AdamW8bit optimizer (Dettmers et al., 2022), a cosine annealing learning rate scheduler, and enabled gradient checkpointing. We also utilized the Liger-Kernel (Hsu et al., 2024) for further training efficiency optimization and enabled bfloat16 mixed-precision training. Other training hyperparameters are provided in Table 10. For SQL generation during inference, we employed greedy decoding.